23 January 2008

RHIC Running:  The problem in the L-10 RF cavity was fixed behind a store yesterday.  A fill was lost yesterday when a collimator (Yellow H2) lost its position and ran into the beam due to a limit switch problem.  After fixing the limit switch, the next fill was lost due to beam losses at cogging.  And the Yellow stochastic cooling kicker in tank 3 failed to insert, so it was left disabled. 

Good fill last night.  APEX began at 5:00AM, except that the experimenters (STAR) could not turn off their high voltage again.   APEX will continue until we start running again at 5:00PM.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020